6. A house wall consists of a 10 cm thick layer of brick (πœ… = 0.68 π‘Šπ‘šβˆ’1𝐾

βˆ’1

), a 15 cmΒ 

thick layer of fiberglass insulation (πœ… = 0.038 π‘Šπ‘šβˆ’1𝐾

βˆ’1

), and a 1 cm thick layer ofΒ 

gypsum board (πœ… = 0.48 π‘Šπ‘šβˆ’1𝐾

βˆ’1

). The interior surface temperature is 22℃ and exteriorΒ 

surface temperature is 5 ℃.

(a) What is the overall R-value of this wall?

(b) What is the heat flux through this wall?


1
Expert's answer
2021-11-18T10:16:51-0500

a)

"R=R_1+R_2+R_3=\frac{\sigma_1}{\kappa_1}+\frac{\sigma_2}{\kappa_2}+\frac{\sigma_3}{\kappa_3}=4.115~\frac{m^2\cdot °C}W,"

b)

"Q=R\Delta T=70~\frac W{m^2}."